OpenStudy (anonymous):

What number replaces the ∆? 5 = 5 5 + 2 = 5 + ∆

OpenStudy (horsegirl27):

Well, do you know what property this is showing?

OpenStudy (horsegirl27):

Actually, we should start with, do you know your math properties?

OpenStudy (horsegirl27):

This is showing commutative property. That means it doesn't matter what order the numbers are written in. ex. a+b=b+a or 1+2=2+1

OpenStudy (horsegirl27):

That should help you figure out what number replaces the ∆ :)

OpenStudy (anonymous):

:)

OpenStudy (horsegirl27):

You figure it out? :P